OpenStudy (anonymous):

What value completes the square for the expression? x^2 + 14x

OpenStudy (p0sitr0n):

+49

OpenStudy (anonymous):

so just 49 is the answer ?

OpenStudy (p0sitr0n):

assuming (x+7)^2

OpenStudy (p0sitr0n):

guess so

OpenStudy (tkhunny):

You must ponder this until you see it. (x+a)^2 = x^2 + 2ax + a^2 You have 2a = 14. You need a^2.